OpenStudy (anonymous):

Given: X = r + 2, Y = 2r - 9, and Z = r² + 17r + 30. Simplify [X · Y - Z] ÷ X. a) -r - 24 b) r - 24 c) 2r2 - 6r - 33 d) 2r2 + 6r + 33

OpenStudy (anonymous):

(B) factor Z=(r+15)(r+2) so, [(X)(Y)-Z] = (r+2)(2r-9)-(r+15)(r+2) take out CF = (r+2)[(2r-9)-(r+15)] now divide the whole thing by X will get = (2r-9)-(r+15) = r-24